Solve and check 2/3 - 3/x = 1/(x-6). Thanks for your help in solving this.

Answer by stanbon(75887)

2/3 - 3/x = 1/(x-6)
--------------------------
2x(x-6) - 3*3(x-6) = 3x
2x^2 - 12x - 9x + 54 = 3x
2x^2 -24x + 54 = 0
x^2 -12x + 27 = 0
x^2 -3x -9x + 27 = 0
x(x-3) - 9(x-3) = 0
(x-3)(x-9) = 0
===================
x = 3 or x = 9
===================
Cheers,
Stan H.